# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 5.2.3 Consistency/Prompt Robustness Benchmarks
  • SCORE (Systematic Consistency and Robustness Evaluation): SCORE (Nalbandyan et al., 2025) provides a framework for non-adversarial robustness evaluation. It assesses models on existing benchmarks (like MMLU-Pro, AGIEval, MATH) under three specific perturbation scenarios :
@@
    • Choice Order Robustness: For MCQ datasets, shuffling the order of answer choices while keeping the correct answer logically in the same position. SCORE reports the range (min/max) of accuracy across these scenarios and the overall Consistency Rate (CR) (Nalbandyan et al., 2025).
  • Paraphrasing Benchmarks: Methodologies that involve generating multiple paraphrased versions of instructions or queries and measuring the consistency of the LLM’s responses (Barbero et al., 2025). This directly tests robustness to linguistic variation.
  • Structural Variation Tests: Evaluating how changes in prompt structure, such as altering the position of the question within the prompt (beginning, middle, end), affect model performance and consistency (Nalbandyan et al., 2025).
@@ 5.2.4 Fairness/Bias Robustness Benchmarks
  • FLEX (Fairness Benchmark in LLM under Extreme Scenarios): FLEX (Jung et al., 2025) focuses on the robustness of fairness. It takes samples from existing fairness benchmarks (like BBQ, CrowS-Pairs, StereoSet) where baseline models typically provide unbiased responses. It then applies adversarial prompts using techniques like persona injection (instructing the LLM to adopt a biased persona), competing objectives (presenting conflicting goals that might induce bias), or text attacks (subtle manipulations of the input) (Jung et al., 2025). FLEX evaluates whether the model maintains its unbiased stance even under these extreme, bias-inducing conditions.
@@ 5.2.5 Task-Specific Robustness Benchmarks
    • ReClor-plus, LogiQA-plus, LogiQAv2-plus (Bao et al., 2025): These extend standard logical reasoning datasets (ReClor, LogiQA, LogiQAv2) by introducing structural variations to the multiple-choice questions. These variations include randomly shuffling the answer options or replacing the correct answer with a “none of the other options is correct” choice, testing if the model relies on superficial patterns or truly understands the logic (Bao et al., 2025).
    • R2ATA (Gan et al., 2024): Assesses reasoning robustness specifically against adversarial typographical errors introduced into the problem statements of reasoning tasks (Gan et al., 2024).
@@
    • CoderEval (Li et al., 2025d): While primarily a code generation benchmark, it has been used in studies focusing on code robustness issues beyond simple functional correctness, such as identifying missing input validation or error handling checks by comparing generated code to human references (Li et al., 2025d).
    • Benchmarks evaluating robustness to perturbations in the natural language instructions provided for code generation tasks are also emerging (Barbero et al., 2025).
@@ 5.2.6 Hallucination/Factuality Benchmarks
  • TruthfulQA: This benchmark (Cecchini et al., 2024) is designed to measure whether LLMs are truthful in generating answers to questions where humans often hold common misconceptions. It tests if models avoid repeating false beliefs prevalent in their training data, distinguishing truthfulness from mere imitation (Yang et al., 2025a). It is also used widely for evaluating factuality robustness with an updated version, TruthfulQA-Enhanced (Yang et al., 2025a).
@@
  • HaluEval: A large-scale benchmark (35,000 samples) focusing on the LLM’s ability to recognize hallucinations (Li et al., 2023). It includes generated and human-annotated samples across QA, knowledge-grounded dialogue, text summarization, and general user queries. Models are evaluated on their ability to correctly classify provided text as containing hallucinations (“Yes”) or not (“No”)
@@
  • FELM (Factuality Evaluation of large Language Models): FELM (Chen et al., 2023b) provides fine-grained, segment-level factuality annotations for LLM-generated responses. For each segment, human annotators provide a factuality label (correct/incorrect), identify the error type if incorrect, and list reference links supporting the judgment. It covers diverse domains including world knowledge, science, math, and reasoning.
@@
  • FACTOR: This benchmark (Muhlgay et al., 2024) automatically transforms a given factual corpus (e.g., Wikipedia, news articles) into a set of evaluation instances. Each instance consists of a context, a factually correct completion, and several plausible but factually incorrect variations (generated based on different error types like entity substitution, negation, and numerical errors). It evaluates an LM by measuring whether it assigns a higher likelihood (probability) to the factually correct completion compared to all the incorrect alternatives in a generation setting (Muhlgay et al., 2024).
@@
  • DefAn: A large dataset ( 75k prompts) across eight domains (including Sports, Census data, Nobel Prize) designed to elicit definitive, concise factual answers (Rahman et al., 2024). It evaluates three aspects: Factual Contradiction Hallucination (FCH rate - % of responses with incorrect facts), Prompt Misalignment Hallucination (PMH rate - % of responses deviating from prompt instructions/format), and Response Consistency (RC - consistency of claims across 15 paraphrased prompts) (Rahman et al., 2024).
@@
  • UMWP (Unanswerable Math Word Problem): This benchmark (Sun et al., 2024a) evaluates hallucination in the context of mathematical reasoning. It consists of math word problems that are intentionally designed to be unanswerable (e.g., due to missing information). It tests whether LLMs correctly identify these problems as unanswerable or if they hallucinate by attempting to provide a numerical solution.
@@
  • Code Hallucination Benchmarks (HalluCode, CodeHalu, Collu-Bench): These benchmarks (Liu et al., 2024a) are specifically designed for the code domain. They often include taxonomies of code-specific hallucinations (e.g., intent conflicting, context inconsistency, dead code, knowledge conflicting (Liu et al., 2024a); mapping, naming, resource, logic hallucinations (Tian et al., 2025)). They evaluate an LLM’s ability to generate correct code and/or recognize hallucinations in generated code snippets. Collu-Bench, for instance, focuses on predicting hallucination locations based on generation log probabilities and execution feedback (Jiang et al., 2024).
@@
  • Multimodal Hallucination Benchmarks (AMBER, POPE, MME, HQHBench, ODE, etc.): These benchmarks evaluate hallucinations in MLLMs, where the generated text might conflict with the provided visual (or other modality) input (Bai et al., 2025b).
@@
    • AMBER (Bai et al., 2025b; Wang et al., 2024b) is notable for being an LLM-free benchmark. It uses carefully annotated images and an automated pipeline to evaluate MLLM hallucinations across both generative and discriminative tasks, covering existence (object presence), attribute (object properties), and relation (spatial/contact relations between objects) hallucinations without relying on another LLM for judgment (Wang et al., 2024b). While other benchmarks like ODE (Tu et al., 2024) propose a dynamic, open-set protocol to generate novel evaluation samples and mitigate data contamination.
@@
The diversity of these hallucination benchmarks highlights the complexity of the phenomenon. Some test the generation of truthful content (TruthfulQA, FACTOR, DefAn), while others test the recognition of flawed content (HaluEval, HalluCode recognition tasks). They cover different domains (general knowledge, dialogue, code, math, multimodal) and rely on different evaluation paradigms (reference-based, reference-free, likelihood-based, classification-based). A comprehensive assessment of an LLM’s factuality and hallucination tendencies requires leveraging multiple benchmarks that cover these different facets (Liu et al., 2024a). Furthermore, the field is actively evolving to address in traditional static benchmarks. There is a noticeable trend towards developing benchmarks that are dynamic (generating test cases on-the-fly to prevent contamination, e.g., ODE, DyVal) (Tu et al., 2024), LLM-free (avoiding reliance on potentially unreliable LLM judges, e.g., AMBER) (Wang et al., 2024b). These advancements aim to improve the reliability and validity of LLM robustness evaluations.
6 Discussion
The topic of robustness of LLMs intersects with several active research areas, leading to a growing body of survey papers and literature reviews. These works provide valuable context and consolidate findings within specific sub-domains or from particular perspectives. Acknowledging and synthesizing these existing reviews is crucial for understanding the broader landscape.
Numerous surveys cover LLMs in general, often touching upon robustness as one aspect among many. A significant number focus on the concept of Trustworthiness in LLMs, consistently positioning robustness as a key pillar alongside fairness, safety, transparency, accountability, privacy, and ethics(Zhao et al., 2025; Aljohani et al., 2025). These trustworthiness surveys emphasize that robustness is essential for achieving reliable and ethically sound deployment, particularly in sensitive applications like healthcare. Some adopt a lifecycle or module-oriented taxonomy (input, model, toolchain, output) to analyze risks, including robustness failures, and discuss mitigations stage-by-stage. Specific dimensions of robustness have also received dedicated survey attention. For instance, adversarial robustness is a frequent topic, with reviews covering attack taxonomies (e.g., character-, word-, sentence-level), defence mechanisms like adversarial training, and evaluation methods. Surveys focusing on Multimodal LLMs (MLLMs) highlight the unique challenges posed by inter-modality dependencies and cross-modal attacks (Jiang et al., 2025). Out-of-Distribution (OOD) Robustness in NLP is another area with dedicated reviews, discussing different types of distribution shifts, benchmark limitations, and evaluation protocols (Lu et al., 2024). The challenge of Shortcut Learning has also been surveyed extensively, identifying it as a fundamental problem underlying poor generalization and adversarial vulnerability in language models (Du et al., 2022). These surveys often categorize types of shortcuts (lexical, syntactic, etc.), discuss their origins in data and training, and review detection and mitigation techniques. Additionally, surveys on Data Augmentation for NLP often discuss its role in improving model robustness. Resources like the Awesome-LLM-Uncertainty-Reliability-Robustness repository also compile relevant papers and categorize them under headings like Invariance, Distribution Shift, OOD, Adversarial, etc. (Ding et al., 2024; Zhang, 2023).
Synthesizing across these diverse surveys reveals several recurring themes and key findings:
  • Centrality of Data Issues: Surveys consistently point to data quality, biases, and artifacts as primary sources of non-robustness, particularly through the mechanism of shortcut learning (Du et al., 2022).
  • Limitations of Standard Training: The inadequacy of standard ERM training for robustness and the challenges introduced by alignment procedures (like RLHF and the associated alignment tax) are common discussion points.
  • Need for Better Evaluation: A widespread consensus exists on the limitations of current evaluation practices and the need for more challenging, standardized, and comprehensive benchmarks and metrics for robustness (Ailem et al., 2024).
  • Diversity of Mitigation: Surveys showcase a wide array of mitigation strategies spanning the entire LLM lifecycle, reinforcing the idea that robustness requires a multi-faceted approach.
The existence of numerous surveys, while indicating intense research activity, also suggests a potentially fragmented landscape where different communities might use varying terminology or focus on specific aspects of the broader robustness problem. For example, issues discussed under “shortcut learning” are directly relevant to “OOD generalization” and “adversarial vulnerability” (Shaw et al., 2025). A key contribution of a comprehensive survey like this one is to bridge these perspectives and synthesize the common underlying principles and challenges. Furthermore, existing surveys often highlight gaps and open questions that remain pertinent. The lack of unified definitions and practical frameworks for operationalising trustworthiness (including robustness) is noted (de Cerqueira et al., 2025). The difficulty in evaluating generative tasks robustly is frequently mentioned. Understanding and mitigating more subtle or complex shortcuts beyond simple lexical cues remains an area for research (Howe et al., 2025; Zhou et al., 2024c; Yuan et al., 2023). Finally, ensuring that robustness improvements scale effectively with increasing model size is an ongoing concern.
@@ 7 Challenges and Future work
The field has made notable progress in identifying and characterizing robustness challenges in LLMs. While awareness of these vulnerabilities is now widespread, and numerous mitigation strategies and evaluation frameworks have emerged, even state-of-the-art LLMs demonstrate brittleness across critical dimensions. Robustness does not consistently correlate with model scale as larger models are not inherently more robust, and scaling trends often show weak or inconsistent patterns (Zhou et al., 2024c; Yang et al., 2024d). Safety-aligned models remain susceptible to sophisticated jailbreaking attempts, while performance may deteriorate sharply with minor distribution shifts or input perturbations (Oh & Demberg, 2025; Yuan et al., 2023). Ultimately, achieving comprehensive and reliable robustness persists as a fundamental challenge for the field.
Major Challenges and Limitations
@@
    Scalability of Defences: Many promising defence mechanisms, particularly those involving adversarial example generation during training (AT) or complex inference-time checks, are computationally expensive and may not scale efficiently to trillion-parameter models (Lee et al., 2024b; Howe et al., 2025). Ensuring robustness improvements keep pace with model scaling is crucial but not guaranteed.
@@
    Robustness-Utility Trade-offs: Aggressively optimizing for robustness can often negatively impact the model’s general capabilities or performance on clean, standard tasks (Lin et al., 2024b; Lee et al., 2024b). The alignment tax associated with RLHF is a prime example of it. Finding the right balance and developing methods that improve robustness without sacrificing utility is a critical ongoing challenge.
@@
    Evaluation Gaps: Current evaluation methodologies have limitations. There is a need for more comprehensive benchmarks covering a wider range of robustness dimensions simultaneously, including more subtle and complex failure modes (e.g., sophisticated shortcuts). Evaluating the robustness of generative capabilities remains particularly difficult compared to classification tasks (Ailem et al., 2024). Ensuring benchmarks are realistic, efficient, and resistant to “teaching to the test” is also vital.
@@
    Theoretical Understanding: Our theoretical understanding of why LLMs exhibit certain robustness failures like shortcut learning, understanding how model architecture (e.g., attention mechanisms, MoE) influences robustness and vulnerabilities or the alignment tax and why certain defences work is still developing (Ye et al., 2024). Deeper theoretical insights could guide the development of more principled and effective mitigation strategies.
@@
Robust Evaluation Methodologies: A critical need exists for more reliable, comprehensive, and standardized benchmarks and evaluation protocols.52 This includes developing robust LLM judges 59, methods for verifying judge outputs 52, and metrics that capture real-world failure modes more effectively.29
Future Research Directions

### v2

@@ 5.2.3 Consistency/Prompt Robustness Benchmarks
  • SCORE (Systematic Consistency and Robustness Evaluation): SCORE (nalbandyan2025scoresystematicconsistencyrobustness) provides a framework for non-adversarial robustness evaluation. It assesses models on existing benchmarks (like MMLU-Pro, AGIEval, MATH) under three specific perturbation scenarios :
@@
    • Choice Order Robustness: For MCQ datasets, shuffling the order of answer choices while keeping the correct answer logically in the same position. SCORE reports the range (min/max) of accuracy across these scenarios and the overall Consistency Rate (CR) (nalbandyan2025scoresystematicconsistencyrobustness).
  • Paraphrasing Benchmarks: Methodologies that involve generating multiple paraphrased versions of instructions or queries and measuring the consistency of the LLM’s responses (barbero2025llmsattendtoken). This directly tests robustness to linguistic variation.
  • Structural Variation Tests: Evaluating how changes in prompt structure, such as altering the position of the question within the prompt (beginning, middle, end), affect model performance and consistency (nalbandyan2025scoresystematicconsistencyrobustness).
@@ 5.2.4 Fairness/Bias Robustness Benchmarks
  • FLEX (Fairness Benchmark in LLM under Extreme Scenarios): FLEX (jung2025flexbenchmarkevaluatingrobustness) focuses on the robustness of fairness. It takes samples from existing fairness benchmarks (like BBQ, CrowS-Pairs, StereoSet) where baseline models typically provide unbiased responses. It then applies adversarial prompts using techniques like persona injection (instructing the LLM to adopt a biased persona), competing objectives (presenting conflicting goals that might induce bias), or text attacks (subtle manipulations of the input) (jung2025flexbenchmarkevaluatingrobustness). FLEX evaluates whether the model maintains its unbiased stance even under these extreme, bias-inducing conditions.
@@ 5.2.5 Task-Specific Robustness Benchmarks
    • ReClor-plus, LogiQA-plus, LogiQAv2-plus (bao2025assessingenhancingrobustnesslarge): These extend standard logical reasoning datasets (ReClor, LogiQA, LogiQAv2) by introducing structural variations to the multiple-choice questions. These variations include randomly shuffling the answer options or replacing the correct answer with a “none of the other options is correct” choice, testing if the model relies on superficial patterns or truly understands the logic (bao2025assessingenhancingrobustnesslarge).
    • R2ATA (gan2024reasoningrobustnessllmsadversarial): Assesses reasoning robustness specifically against adversarial typographical errors introduced into the problem statements of reasoning tasks (gan2024reasoningrobustnessllmsadversarial).
@@
    • CoderEval (unknowntask): While primarily a code generation benchmark, it has been used in studies focusing on code robustness issues beyond simple functional correctness, such as identifying missing input validation or error handling checks by comparing generated code to human references (unknowntask).
    • Benchmarks evaluating robustness to perturbations in the natural language instructions provided for code generation tasks are also emerging (barbero2025llmsattendtoken).
@@ 5.2.6 Hallucination/Factuality Benchmarks
  • TruthfulQA: This benchmark (cecchini-etal-2024-holistic) is designed to measure whether LLMs are truthful in generating answers to questions where humans often hold common misconceptions. It tests if models avoid repeating false beliefs prevalent in their training data, distinguishing truthfulness from mere imitation (yang2025hallucinationdetectionlargelanguage). It is also used widely for evaluating factuality robustness with an updated version, TruthfulQA-Enhanced (yang2025hallucinationdetectionlargelanguage).
@@
  • HaluEval: A large-scale benchmark (35,000 samples) focusing on the LLM’s ability to recognize hallucinations (li2023haluevallargescalehallucinationevaluation). It includes generated and human-annotated samples across QA, knowledge-grounded dialogue, text summarization, and general user queries. Models are evaluated on their ability to correctly classify provided text as containing hallucinations (“Yes”) or not (“No”)
@@
  • FELM (Factuality Evaluation of LLMs): FELM (chen2023felm) provides fine-grained, segment-level factuality annotations for LLM-generated responses. For each segment, human annotators provide a factuality label (correct/incorrect), identify the error type if incorrect, and list reference links supporting the judgment. It covers diverse domains including world knowledge, science, math, and reasoning.
@@
  • FACTOR: This benchmark (muhlgay2024generatingbenchmarksfactualityevaluation) automatically transforms a given factual corpus (e.g., Wikipedia, news articles) into a set of evaluation instances. Each instance consists of a context, a factually correct completion, and several plausible but factually incorrect variations (generated based on different error types like entity substitution, negation, and numerical errors). It evaluates an LM by measuring whether it assigns a higher likelihood (probability) to the factually correct completion compared to all the incorrect alternatives in a generation setting (muhlgay2024generatingbenchmarksfactualityevaluation).
@@
  • DefAn: A large dataset ( 75k prompts) across eight domains (including Sports, Census data, Nobel Prize) designed to elicit definitive, concise factual answers (rahman2024defandefinitiveanswerdataset). It evaluates three aspects: Factual Contradiction Hallucination (FCH rate - % of responses with incorrect facts), Prompt Misalignment Hallucination (PMH rate - % of responses deviating from prompt instructions/format), and Response Consistency (RC - consistency of claims across 15 paraphrased prompts) (rahman2024defandefinitiveanswerdataset).
@@
  • UMWP (Unanswerable Math Word Problem): This benchmark (sun2024benchmarkinghallucinationlargelanguage) evaluates hallucination in the context of mathematical reasoning. It consists of math word problems that are intentionally designed to be unanswerable (e.g., due to missing information). It tests whether LLMs correctly identify these problems as unanswerable or if they hallucinate by attempting to provide a numerical solution.
@@
  • Code Hallucination Benchmarks (HalluCode, CodeHalu, Collu-Bench): These benchmarks (liu2024exploringevaluatinghallucinationsllmpowered) are specifically designed for the code domain. They often include taxonomies of code-specific hallucinations (e.g., intent conflicting, context inconsistency, dead code, knowledge conflicting (liu2024exploringevaluatinghallucinationsllmpowered); mapping, naming, resource, logic hallucinations (tian2025codehaluinvestigatingcodehallucinations)). They evaluate an LLM’s ability to generate correct code and/or recognize hallucinations in generated code snippets. Collu-Bench, for instance, focuses on predicting hallucination locations based on generation log probabilities and execution feedback (jiang2024collubenchbenchmarkpredictinglanguage).
@@
  • Multimodal Hallucination Benchmarks (AMBER, POPE, MME, HQHBench, ODE, etc.): These benchmarks evaluate hallucinations in MLLMs, where the generated text might conflict with the provided visual (or other modality) input (bai2025hallucinationmultimodallargelanguage).
@@
    • AMBER (bai2025hallucinationmultimodallargelanguage; wang2024amberllmfreemultidimensionalbenchmark) is notable for being an LLM-free benchmark. It uses carefully annotated images and an automated pipeline to evaluate MLLM hallucinations across both generative and discriminative tasks, covering existence (object presence), attribute (object properties), and relation (spatial/contact relations between objects) hallucinations without relying on another LLM for judgment (wang2024amberllmfreemultidimensionalbenchmark). While other benchmarks like ODE (tu2024odeopensetevaluationhallucinations) propose a dynamic, open-set protocol to generate novel evaluation samples and mitigate data contamination.
@@
The diversity of these hallucination benchmarks highlights the complexity of the phenomenon. Some test the generation of truthful content (TruthfulQA, FACTOR, DefAn), while others test the recognition of flawed content (HaluEval, HalluCode recognition tasks). They cover different domains (general knowledge, dialogue, code, math, multimodal) and rely on different evaluation paradigms (reference-based, reference-free, likelihood-based, classification-based). A comprehensive assessment of an LLM’s factuality and hallucination tendencies requires leveraging multiple benchmarks that cover these different facets (liu2024exploringevaluatinghallucinationsllmpowered). Furthermore, the field is actively evolving to address in traditional static benchmarks. There is a noticeable trend towards developing benchmarks that are dynamic (generating test cases on-the-fly to prevent contamination, e.g., ODE, DyVal) (tu2024odeopensetevaluationhallucinations), LLM-free (avoiding reliance on potentially unreliable LLM judges, e.g., AMBER) (wang2024amberllmfreemultidimensionalbenchmark). These advancements aim to improve the reliability and validity of LLM robustness evaluations.
@@ 6 Challenges and Future Work
The field has made notable progress in identifying and characterizing robustness challenges in LLMs. While awareness of these vulnerabilities is now widespread, and numerous mitigation strategies and evaluation frameworks have emerged, even state-of-the-art LLMs demonstrate brittleness across critical dimensions. Robustness does not consistently correlate with model scale as larger models are not inherently more robust, and scaling trends often show weak or inconsistent patterns (zhou2024navigatingshortcutmazecomprehensive; yang2024assessingadversarialrobustnesslarge). Safety-aligned models remain susceptible to sophisticated jailbreaking attempts, while performance may deteriorate sharply with minor distribution shifts or input perturbations (safety; Oh_Demberg_2025; yuan2023revisiting). Ultimately, achieving comprehensive and reliable robustness persists as a fundamental challenge for the field.
6.1 Major Challenges and Limitations
@@
    Scalability of Defences: Many promising defence mechanisms, particularly those involving adversarial example generation during training (AT) or complex inference-time checks, are computationally expensive and may not scale efficiently to trillion-parameter models (Lee2024HarmAugED; howe2025scalingtrendslanguagemodel). Ensuring robustness improvements keep pace with model scaling is crucial but not guaranteed.
@@
    Robustness-Utility Trade-offs: Aggressively optimizing for robustness can often negatively impact the model’s general capabilities or performance on clean, standard tasks (lin2024mitigatingalignmenttaxrlhf; Lee2024HarmAugED). The alignment tax associated with RLHF is a prime example of it. Finding the right balance and developing methods that improve robustness without sacrificing utility is a critical ongoing challenge.
@@
    Evaluation Gaps: Current evaluation methodologies have limitations. There is a need for more comprehensive benchmarks covering a wider range of robustness dimensions simultaneously, including more subtle and complex failure modes (e.g., sophisticated shortcuts). Evaluating the robustness of generative capabilities remains particularly difficult compared to classification tasks (ailem2024examiningrobustnessllmevaluation). Ensuring benchmarks are realistic, efficient, and resistant to “teaching to the test” is also vital.
@@
    Theoretical Understanding: Our theoretical understanding of why LLMs exhibit certain robustness failures like shortcut learning, understanding how model architecture (e.g., attention mechanisms, MoE) influences robustness and vulnerabilities or the alignment tax and why certain defences work is still developing (ye2024spuriouscorrelationsmachinelearning). Deeper theoretical insights could guide the development of more principled and effective mitigation strategies.
@@
  6. 6.
    Robust Evaluation Methodologies: A critical need exists for more reliable, comprehensive, and standardized benchmarks and evaluation protocols. This includes developing robust LLM judges, methods for verifying judge outputs, and metrics that capture real-world failure modes more effectively.
6.2 Future Research Directions